A group plans to share the cost equally for a $200,000 plane,the group wants to find 5 more people to join ,so that the cost will decrease $2,000 per person .How people are currently in the group.

Found 2 solutions by mananth, Theo:
Answer by mananth(16946)

let the people in the group be x
cost / person = 200000/x
increase by 5
x+5
200000/x+5
200000 /x -2000 = 200000/x+5
200000 -2000x /x = 200000 / x+5
200000x -10000x +1000000 = 200000x
-10000x=-1000000
Number of people now
x=100



Answer by Theo(13342)

cost of the plane is $200,000

There are x number of people in the group.

If the group adds 5 people, there will be x+5 people in the group and the cost per person will be $2,000

The cost per person would be the total cost divided by the number of people in the group.

Your equation would be:

$200,000/(x+5) = $2,000

Multiply both sides of this equaqtion by (x+5) to get:

$200,000 = $2,000 * (x+5)

Divide both sides of this equation by $2,000 to get:

100 = x + 5

Subtract 5 from each side of this equation to get:

x = 95

There are 95 people originally in this group.

They add 5 and the number of people in the group becomes 100.

$200,000 / 100 = $2,000 per person.